OpenStudy (logo):

A baseball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t^2 + 32t + 10. h is height, t is time in seconds. How many seconds after the ball is thrown will it reach its maximum and minimum height? Need help!

OpenStudy (logo):

its 9 not 10 for the equation -16t^2 + 32t + 9

OpenStudy (welshfella):

convert to vertex form:- -16t^2 + 32t + 9 = -16[ t^- 2t] + 9 = -16[ (t - 1)^2 - 1] + 9 = -16(t - 1)^2 + 25 the maximum height is 25 after 1 second ( when t = 1)

OpenStudy (welshfella):

it will take the same time to reach the ground as it took to reach maximum so minimum time = ?

OpenStudy (logo):

minimum time = 1 ? @welshfella

OpenStudy (logo):

@welshfella oh wait I'm sorry, it only asked for its maximum height not the minimum. I was so confused about the minimum. I even tried it this way by finding the vertex first using t = -b/2a and got 25 feet in the end.